OpenStudy (anonymous):

how do you rewrite the expression in simplified expontential form sqrt x^3* ^3sqrt x^2

OpenStudy (amistre64):

x^(3/2) * x^(2/3) = x^(3/2 + 2/3) = x^(13/6)

OpenStudy (amistre64):

x^2 * x^(1/6) maybe?